How to divide value 1180 into ratio1/5:1/6:1/8
Kindly suggest the method also

Answer by Alan3354(69443)

How to divide value 1180 into ratio1/5:1/6:1/8
--------------------
1/5 + 1/6 + 1/8 = (24 + 20 + 15)/120 = 59/120
1180/59 = 20
---
--> 24*20:20*20:15*20
= 480:400:300